# MeanSparse: Post-Training Robustness Enhancement Through Mean-Centered Feature Sparsification

### v2

@@ A.4 Adaptive Attack
where $\mu_{\mathrm{ch}}$ and $\sigma_{\mathrm{ch}}$ are the statistics of the MeanSparse module, and $K$ represents the representation dimension. Thus, whenever a feature in the input of MeanSparse is affected by this module for the input image $\bm{x}$, the corresponding mask value becomes $1$; otherwise, it is 00. As a result, the second term in the objective function of the adaptive PGD attack equation 9 penalizes perturbations that lead to changes in the features affected by MeanSparse. In the ideal case, the perturbation is designed such that none of the MeanSparse modules can block its effect in their input (we know that the changes in the input of the $i$-th MeanSparse module are $\bm{z}_{i}-\bm{z}_{i}^{d}$). Thus, the attacker can bypass the MeanSparse, which we denote as the Adaptive-$(p,q)$ attack.
To test the adaptive attack, we solve the optimization problem equation 9 using gradient ascent with a step size of $\mu=\frac{1}{255}$, a number of iterations $N=50$, and $\epsilon=\frac{8}{255}$. When $\lambda=0$, problem equation 9 reduces to the well-known PGD-50 attack, while for larger values of $\lambda$, we obtain an adaptive attack. We conducted experiments using this adaptive attack on a ResNet-18 model with a GELU activation function, which was adversarially trained on the CIFAR-10 dataset. Table 5 compares the results of the PGD-50 attack with the Adaptive-$(1,1)$ and Adaptive-$(2,2)$ attacks before (Base) and after integration ($\alpha=0.20$ and $\alpha=0.35$) with MeanSparse. For the base model, all the attacks are equivalent, as $\bm{m}_{i}$ is a zero matrix for all MeanSparse modules. As $\alpha$ increases, the accuracy after applying our adaptive attack decreases compared to the PGD-50 attack, demonstrating the effectiveness of our adaptation. While the adapted attack is more powerful than the original PGD-50 attack, the improvement gap in robustness resulting from MeanSparse is still significant. For the base model, the robust accuracy is $51.56\%$, while the worst accuracies are $53.56\%$ and $55.26\%$ for $\alpha=0.20$ and $\alpha=0.35$, respectively, indicating a clear improvement in robustness, while the clean accuracy remains almost unchanged.
Table 5: Clean, PGD-50 and Adaptive Attack accuracy for a ResNet-18 model with GELU activation over CIFAR-10 dataset before (Base) and after integration with MeanSparse
@@
A.5 Adversarial Training
MeanSparse is designed to integrate with models trained using adversarial training. In previous experiments, PGD [29] was used for adversarial training. In this section, we compare the PGD results with those of TRADES [54] for adversarial training. For TRADES adversarial training, we set $\beta=0.6$. For the other properties of adversarial training, we adopt the parameters specified in Section 2. Table 6 compares the results for base adversarially trained models and models integrated with MeanSparse. For TRADES, MeanSparse can still improve the APGD accuracy while maintaining clean accuracy. This table demonstrates overlapping information gathered around the mean feature values for both PGD and TRADES as adversarial training methods. In both methods, the information around the feature mean with a value of $\alpha\simeq 0.2$ is almost uninformative for clean data but can be utilized by attackers. Thus, the performance of MeanSparse represents generalizability across both PGD and TRADES adversarial training methods.
Table 6: Clean and APGD accuracy of ResNet-18 model with GELU activation function over CIFAR-10 test set before (Base) and after integration with MeanSparse technique for different value of $\alpha$ and adversarial training based on PGD [29] and TRADES [54]
Adversarial training
@@ A.5 Adversarial Training
A.6 Standard Training
In the previous experiments, we explore the effect of two different adversarial training methods on performance improvement by MeanSparse. In this experiment, we test MeanSparse integration over a standard trained model. Table 7 represents the APGD accuracy over a ResNet-18 model with standard training for $\epsilon$ values starting from $1/255$ to $8/255$. The first column ($\alpha=0$) is the base model performance while the next two columns ($\alpha=0.2$ and $\alpha=0.35$) represent the models integrated with MeanSparse. In all cases, the attacker can fool the model with imperceptible perturbation ($\epsilon\leq 4/255$) which necessitates the integration of MeanSparse with adversarially trained models.
Table 7: APGD accuracy of standard trained ResNet-18 model with GELU activation function over CIFAR-10 test set before (Base denoted by $\alpha=0$ in the table) and after the integration of MeanSparse technique for different attack power $\epsilon$ and two different $\alpha$ values ($0.20$ and $0.35$). Clean accuracy is provided next to each $\alpha$ value.
$\alpha$
@@ A.6 Standard Training
A.7 Feature Centerization Type
The MeanSparse technique applies sparsification over mean-centered features across each channel in the input tensor to the activation function. In this section, we inspect the importance of the reference for feature centralization before sparsification in our proposed technique. For this purpose, we compare three cases where the reference for centering features is the channel mean ($\bm{\mu}_{ch}$), zero, and the mean over all features in all channels ($\mu$). For a fair comparison across all cases, we select the sparsification threshold equal to $\alpha\times\sigma_{ch}$, where $\sigma_{ch}$ emphasizes the channel-wise calculation standard deviation. Table 8 represents the result for different options of reference. When using zero as the reference for feature centralization, the improvement in APGD is higher than using the mean as the reference, but on the other hand, the utilization of the model also decreases. One challenging problem is the high rate of clean accuracy reduction when using zero as the reference, which may cause a huge reduction if not selected properly. When using the mean of all channels as the reference, the results in terms of both clean and APGD accuracy metrics underperform compared to the case where the channel mean is used as the reference. The main takeaway from the table is the importance of using channel-wise statistics to maintain utility while improving robustness.
Table 8: Clean and APGD accuracy of ResNet-18 model with GELU activation function over CIFAR-10 test set before (Base) and after the application of MeanSparse technique for different references for feature centralization
Reference for Centralization
zero $\mu$ $\bm{\mu}_{ch}$
@@ A.7 Feature Centerization Type
$Base$ $\mathbf{84.59}$ $49.12$ $\mathbf{84.59}$ $49.12$ $84.59$ $49.12$
@@
A.8 Attack Power
In the previous experiments, the attack power was set to a fixed value of $8/255$ for $\ell_{\infty}$ attack threat which is generally the threshold of imperceptibility [10]. In this experiment, we explore the effectiveness of the MeanSparse technique for a wide range of attack powers. Table 9 represents the APGD accuracy for $\epsilon$ values starting from $1/255$ to $16/255$. For $\alpha=0.2$, the accuracy is smaller than the base model ($\alpha=0$) for $1/255$ and $2/255$ values of $\epsilon$, which is generally due to the reduction of clean accuracy after applying MeanSparse (clean accuracy reduces to $84.32$ from $84.59$). For $\epsilon$ values larger than $2/255$, applying the MeanSparse technique persistently improves the APGD accuracy. The same trend is also observed for $\alpha=0.35$ except for the fact that after $\epsilon=4/255$, we have persistent improvements over the base model.
Table 9: APGD accuracy of ResNet-18 model with GELU activation function over CIFAR-10 test set before (Base denoted by $\alpha=0$ in the table) and after the application of MeanSparse technique for different attack power $\epsilon$ and two different $\alpha$ values ($0.20$ and $0.35$). Clean accuracy is provided next to each $\alpha$ value.
$\alpha$
@@ A.8 Attack Power
A.9 Applying to Specific Activation Functions
@@
Table 10 represents the results for both Single Activation and Cumulative scenarios. One shared remark is that the starting layers are more effective in improving robustness. We can attribute this to the abstraction level in different layers of the ResNet-18 model. In the case of Single Activation, when we apply the MeanSparse technique to larger indices, we observe that not only does the robust accuracy decrease, but the clean accuracy increases. This indicates that for even better performance, we can use different $\alpha$ values for activation functions based on their depth level. Note that for the current version of the MeanSparse technique, we use a shared $\alpha$ for all activation functions.
Another aspect worth inspecting in this experiment is how the application of the MeanSparse technique to activation functions in the main path and after the addition of residual data to the main path result affects robustness in terms of APGD accuracy. So we consider two cases in Table 11. In the first case, MeanSparse is applied to all the activation functions in the main path of the ResNet block, denoted by Main Path. In the second case, MeanSparse is applied to all the activation functions after the addition of residual data to the output of the main path, denoted by After Addition. This table reveals an important result. While the application of MeanSparse to both types of activation functions contributes to the robustness of the model, the application of MeanSparse to the main flow of the model (After Addition) is more responsible for the improvements in the ranges where clean accuracy is almost unchanged ($\alpha\simeq 0.2$).
Table 10: Clean and APGD accuracy of ResNet-18 model with GELU activation function over CIFAR-10 test set after the application of MeanSparse technique for Single Activation and Cumulative cases. Single Activation represents the results when MeanSparse is applied to only the index activation function, and Cumulative Activation represents the results when MeanSparse is applied to the index activation function and all previous activation functions (for the based model the Clean and APGD accuracy are $84.59\%$ and $49.12\%$, respectively).
@@
$1$ $84.56$ $49.19$ $84.44$ $49.09$ $\mathbf{84.53}$ $49.38$ $\mathbf{84.31}$ $49.35$
$2$ $84.46$ $49.21$ $84.45$ $49.20$ $84.49$ $49.37$ $83.98$ $49.29$
$3$ $84.50$ $49.15$ $84.33$ $49.19$ $84.45$ $49.52$ $83.72$ $49.69$
$4$ $84.48$ $49.21$ $84.26$ $\mathbf{49.26}$ $84.41$ $49.67$ $83.47$ $50.02$
$5$ $84.54$ $49.24$ $84.36$ $49.25$ $84.38$ $49.86$ $83.14$ $50.29$
$6$ $84.60$ $49.11$ $84.58$ $49.21$ $84.34$ $49.90$ $83.21$ $50.50$
$7$ $84.51$ $49.13$ $84.45$ $49.16$ $84.32$ $49.99$ $83.02$ $50.65$
$8$ $84.58$ $49.19$ $84.59$ $49.21$ $84.34$ $50.12$ $83.05$ $50.86$
$9$ $84.60$ $49.12$ $84.53$ $49.02$ $84.35$ $50.32$ $82.96$ $51.09$
@@
Table 11: Clean and APGD accuracy of ResNet-18 model with GELU activation function over CIFAR-10 test set before (Base denoted by $\alpha=0$ in the table) and after the application of MeanSparse technique for Main Path (application of MeanSparse to activation functions in the main path of residual block) and After Addition (application of MeanSparse to activation functions operating after addition of main path result and residual data) cases.
Layer type
@@
A.10 Attack Type
Figure LABEL:fig:spaf demonstrates the improvement in robustness of leading models in RobustBench in terms of AutoAttack accuracy. In this experiment, we use PGD-50 [29] for robustness measurement. Table 12 represents the Clean and PGD-50 accuracy values for different models before (denoted by Base) and after integration with MeanSparse (denoted by MeanSparse) over CIFAR-10, CIFAR-100 and ImageNet datasets. The results again confirm the improvements in Figure LABEL:fig:spaf as t integration with MeanSparse increases the PGD-50 accuracy in all models.
Table 12: Clean and PGD-50 [29] accuracy of various RobustBench-ranked models [9], both before (Base) and after integration with the MeanSparse technique (models tested here includes WideResNet-94-16 [3] and RaWideResNet-70-16 [34] over CIFAR-10 dataset, WideResNet-70-16 [48] over CIFAR-100 dataset and Swin-L [27] and ConvNeXt-L [26] over ImageNet dataset
CIFAR-10 CIFAR-100 ImageNet
Clean PGD-50 Clean PGD-50 Clean PGD-50 Clean PGD-50 Clean PGD-50

### v3

@@ IV-E Adaptive Attack
where $\mu_{\mathrm{ch}}$ and $\sigma_{\mathrm{ch}}$ are the statistics of the MeanSparse module, and $K$ represents the representation dimension. Thus, whenever a feature in the input of MeanSparse is affected by this module for the input image $\boldsymbol{x}$, the corresponding mask value becomes $1$; otherwise, it is 0. As a result, the second term in the objective function of the adaptive PGD attack (8) penalizes perturbations that lead to changes in the features affected by MeanSparse. In the ideal case, the perturbation is designed such that none of the MeanSparse modules can block its effect in their input (we know that the changes in the input of the $i$-th MeanSparse module are $\boldsymbol{z}_{i}-\boldsymbol{z}_{i}^{d}$). Thus, the attacker can bypass the MeanSparse, which we denote as the Adaptive-$(p,q)$ attack.
To test the adaptive attack, we solve the optimization problem (8) using gradient ascent with a step size of $\mu=\frac{1}{255}$, a number of iterations $N=50$, and $\epsilon=\frac{8}{255}$. When $\lambda=0$, problem (8) reduces to the well-known PGD-50 attack, while for larger values of $\lambda$, we obtain an adaptive attack. We conducted experiments using this adaptive attack on a ResNet-18 model with a GELU activation function, which was adversarially trained on the CIFAR-10 dataset. Table V compares the results of the PGD-50 attack with the Adaptive-$(1,1)$ and Adaptive-$(2,2)$ attacks before (Base) and after integration ($\alpha=0.20$ and $\alpha=0.35$) with MeanSparse. For the base model, all the attacks are equivalent, as $\boldsymbol{m}_{i}$ is a zero matrix for all MeanSparse modules. As $\alpha$ increases, the accuracy after applying our adaptive attack decreases compared to the PGD-50 attack, demonstrating the effectiveness of our adaptation. While the adapted attack is more powerful than the original PGD-50 attack, the improvement gap in robustness resulting from MeanSparse is still significant. For the base model, the robust accuracy is $51.56\%$, while the worst accuracies are $53.56\%$ and $55.26\%$ for $\alpha=0.20$ and $\alpha=0.35$, respectively, indicating a clear improvement in robustness, while the clean accuracy remains almost unchanged.
TABLE V: Clean, PGD-50 and Adaptive Attack accuracy for a ResNet-18 model with GELU activation over CIFAR-10 dataset before (Base) and after integration with MeanSparse
@@
IV-F Adversarial Training
MeanSparse is designed to integrate with models trained using adversarial training. In previous experiments, PGD [1] was used for adversarial training. In this section, we compare the PGD results with those of TRADES [55] for adversarial training. For TRADES adversarial training, we set $\beta=0.6$. For the other properties of adversarial training, we adopt the parameters specified in Section II. Table VI compares the results for base adversarially trained models and models integrated with MeanSparse. For TRADES, MeanSparse can still improve the APGD accuracy while maintaining clean accuracy. This table demonstrates overlapping information gathered around the mean feature values for both PGD and TRADES as adversarial training methods. In both methods, the information around the feature mean with a value of $\alpha\simeq 0.2$ is almost uninformative for clean data but can be utilized by attackers. Thus, the performance of MeanSparse represents generalizability across both PGD and TRADES adversarial training methods.
TABLE VI: Clean and APGD accuracy of ResNet-18 model with GELU activation function over CIFAR-10 test set before (Base) and after integration with MeanSparse technique for different value of $\alpha$ and adversarial training based on PGD [1] and TRADES [55]
@@ IV-F Adversarial Training
IV-G Standard Training
In the previous experiments, we explore the effect of two different adversarial training methods on performance improvement by MeanSparse. In this experiment, we test MeanSparse integration over a standard trained model. Table VII represents the APGD accuracy over a ResNet-18 model with standard training for $\epsilon$ values starting from $1/255$ to $8/255$. The first column ($\alpha=0$) is the base model performance while the next two columns ($\alpha=0.2$ and $\alpha=0.35$) represent the models integrated with MeanSparse. In all cases, the attacker can fool the model with imperceptible perturbation ($\epsilon\leq 4/255$) which necessitates the integration of MeanSparse with adversarially trained models.
TABLE VII: APGD accuracy of standard trained ResNet-18 model with GELU activation function over CIFAR-10 test set before (Base denoted by $\alpha=0$ in the table) and after the integration of MeanSparse technique for different attack power $\epsilon$ and two different $\alpha$ values ($0.20$ and $0.35$). Clean accuracy is provided next to each $\alpha$ value.
@@ IV-G Standard Training
IV-H Feature Centerization Type
The MeanSparse technique applies sparsification over mean-centered features across each channel in the input tensor to the activation function. In this section, we inspect the importance of the reference for feature centralization before sparsification in our proposed technique. For this purpose, we compare three cases where the reference for centering features is the channel mean ($\boldsymbol{\mu}_{ch}$), zero, and the mean over all features in all channels ($\mu$). For a fair comparison across all cases, we select the sparsification threshold equal to $\alpha\times\sigma_{ch}$, where $\sigma_{ch}$ emphasizes the channel-wise calculation standard deviation. Table VIII represents the result for different options of reference. When using zero as the reference for feature centralization, the improvement in APGD is higher than using the mean as the reference, but on the other hand, the utilization of the model also decreases. One challenging problem is the high rate of clean accuracy reduction when using zero as the reference, which may cause a huge reduction if not selected properly. When using the mean of all channels as the reference, the results in terms of both clean and APGD accuracy metrics underperform compared to the case where the channel mean is used as the reference. The main takeaway from the table is the importance of using channel-wise statistics to maintain utility while improving robustness.
TABLE VIII: Clean and APGD accuracy of ResNet-18 model with GELU activation function over CIFAR-10 test set before (Base) and after the application of MeanSparse technique for different references for feature centralization
zero $\mu$ $\boldsymbol{\mu}_{ch}$
@@ IV-H Feature Centerization Type
BaseBase $\mathbf{84.59}$ $49.12$ $\mathbf{84.59}$ $49.12$ $84.59$ $49.12$
@@
IV-I Attack Power
In the previous experiments, the attack power was set to a fixed value of $8/255$ for $\ell_{\infty}$ attack threat which is generally the threshold of imperceptibility [53]. In this experiment, we explore the effectiveness of the MeanSparse technique for a wide range of attack powers. Table IX represents the APGD accuracy for $\epsilon$ values starting from $1/255$ to $16/255$. For $\alpha=0.2$, the accuracy is smaller than the base model ($\alpha=0$) for $1/255$ and $2/255$ values of $\epsilon$, which is generally due to the reduction of clean accuracy after applying MeanSparse (clean accuracy reduces to $84.32$ from $84.59$). For $\epsilon$ values larger than $2/255$, applying the MeanSparse technique persistently improves the APGD accuracy. The same trend is also observed for $\alpha=0.35$ except for the fact that after $\epsilon=4/255$, we have persistent improvements over the base model.
TABLE IX: APGD accuracy of ResNet-18 model with GELU activation function over CIFAR-10 test set before (Base denoted by $\alpha=0$ in the table) and after the application of MeanSparse technique for different attack power $\epsilon$ and two different $\alpha$ values ($0.20$ and $0.35$). Clean accuracy is provided next to each $\alpha$ value.
@@ IV-I Attack Power
IV-J Applying to Specific Activation Functions
@@
Table X represents the results for both Single Activation and Cumulative scenarios. One shared remark is that the starting layers are more effective in improving robustness. We can attribute this to the abstraction level in different layers of the ResNet-18 model. In the case of Single Activation, when we apply the MeanSparse technique to larger indices, we observe that not only does the robust accuracy decrease, but the clean accuracy increases. This indicates that for even better performance, we can use different $\alpha$ values for activation functions based on their depth level. Note that for the current version of the MeanSparse technique, we use a shared $\alpha$ for all activation functions.
Another aspect worth inspecting in this experiment is how the application of the MeanSparse technique to activation functions in the main path and after the addition of residual data to the main path result affects robustness in terms of APGD accuracy. So we consider two cases in Table XIII. In the first case, MeanSparse is applied to all the activation functions in the main path of the ResNet block, denoted by Main Path. In the second case, MeanSparse is applied to all the activation functions after the addition of residual data to the output of the main path, denoted by After Addition. This table reveals an important result. While the application of MeanSparse to both types of activation functions contributes to the robustness of the model, the application of MeanSparse to the main flow of the model (After Addition) is more responsible for the improvements in the ranges where clean accuracy is almost unchanged ($\alpha\simeq 0.2$).
TABLE X: Clean and APGD accuracy of ResNet-18 model with GELU activation function over CIFAR-10 test set after the application of MeanSparse technique for Single Activation and Cumulative cases. Single Activation represents the results when MeanSparse is applied to only the index activation function, and Cumulative Activation represents the results when MeanSparse is applied to the index activation function and all previous activation functions (for the based model the Clean and APGD accuracy are $84.59\%$ and $49.12\%$, respectively).
@@
11 $84.56$ $49.19$ $84.44$ $49.09$ $\mathbf{84.53}$ $49.38$ $\mathbf{84.31}$ $49.35$
22 $84.46$ $49.21$ $84.45$ $49.20$ $84.49$ $49.37$ $83.98$ $49.29$
33 $84.50$ $49.15$ $84.33$ $49.19$ $84.45$ $49.52$ $83.72$ $49.69$
44 $84.48$ $49.21$ $84.26$ $\mathbf{49.26}$ $84.41$ $49.67$ $83.47$ $50.02$
55 $84.54$ $49.24$ $84.36$ $49.25$ $84.38$ $49.86$ $83.14$ $50.29$
66 $84.60$ $49.11$ $84.58$ $49.21$ $84.34$ $49.90$ $83.21$ $50.50$
77 $84.51$ $49.13$ $84.45$ $49.16$ $84.32$ $49.99$ $83.02$ $50.65$
88 $84.58$ $49.19$ $84.59$ $49.21$ $84.34$ $50.12$ $83.05$ $50.86$
99 $84.60$ $49.12$ $84.53$ $49.02$ $84.35$ $50.32$ $82.96$ $51.09$
@@
TABLE XI: Clean and PGD-50 [1] accuracy of various RobustBench-ranked models [16], both before (Base) and after integration with the MeanSparse technique (models tested here includes WideResNet-94-16 [49] and RaWideResNet-70-16 [15] over CIFAR-10 dataset, WideResNet-70-16 [50] over CIFAR-100 dataset and Swin-L [51] and ConvNeXt-L [52] over ImageNet
CIFAR-10 CIFAR-100 ImageNet
Clean PGD-50 Clean PGD-50 Clean PGD-50 Clean PGD-50 Clean PGD-50
TABLE XII: Relative MAE between subset and full-dataset feature statistics for various robust models.
WRN-70-16 (CIFAR-10, $\ell_{2}$) [50] WRN-94-16 (CIFAR-10, $\ell_{\infty}$) [49] WRN-70-16 (CIFAR-100) [50] Swin-L (ImageNet) [51]
Subset (%) Mean Std Mean Std Mean Std Mean Std
10 0.1476 0.0033 0.1286 0.0038 0.3481 0.0044 0.0228 0.0275
25 0.0471 0.0015 0.0707 0.0017 0.0930 0.0019 0.0119 0.0274
50 0.0382 0.0010 0.0520 0.0010 0.1000 0.0010 0.0037 0.0272
75 0.0228 0.0005 0.0205 0.0006 0.0784 0.0006 0.0046 0.0272
TABLE XIII: Clean and APGD accuracy of ResNet-18 model with GELU activation function over CIFAR-10 test set before (Base denoted by $\alpha=0$ in the table) and after the application of MeanSparse technique for Main Path (application of MeanSparse to activation functions in the main path of residual block) and After Addition (application of MeanSparse to activation functions operating after addition of main path result and residual data) cases.
@@
IV-K Attack Type
Figure 3 demonstrates the improvement in robustness of leading models in RobustBench in terms of AutoAttack accuracy. In this experiment, we use PGD-50 [1] for robustness measurement. Table XI represents the Clean and PGD-50 accuracy values for different models before (denoted by Base) and after integration with MeanSparse (denoted by MeanSparse) over CIFAR-10, CIFAR-100 and ImageNet datasets. The results again confirm the improvements in Figure 3 as t integration with MeanSparse increases the PGD-50 accuracy in all models.
IV-L Robustness with Subset-Based Statistics
A practical advantage of MeanSparse lies in its minimal data requirements at post-training. Instead of accessing the full training dataset, the method only requires feature statistics (mean and variance), which can be reliably estimated from subsets of the data. Table XII presents the approximation quality when statistics are computed from 10%, 25%, 50%, and 75% of the training set, evaluated by Relative Mean Absolute Error (MAE). We observe that Relative MAE values below 10% already yield accurate estimates, while values below 5% indicate close alignment with full-data statistics. For instance, with only 25% of CIFAR-10 data, the WideResNet-70-16 model achieves a Relative MAE of approximately 5%, leading to natural and robust accuracies of 95.49% and 87.31%, respectively. These results closely match the full-data performance of 95.49% and 87.28%, demonstrating that MeanSparse maintains reliability even when computed with partial datasets. This finding highlights the scalability of the method to scenarios where full training data access is restricted due to privacy, storage, or efficiency constraints.
V Limitations
The MeanSparse technique proposed in this paper is a post-training method that is useful on top of adversarially trained models. Using MeanSparse for models trained in a standard way does not lead to robustness improvement as we expect based on our intuition introduced in this paper. Additionally, while integrating a model with MeanSparse shows clear improvements against both white-box and black-box non-adaptive attacks, white-box attacks that ignore the MeanSparse transformation during backpropagation and use the identity function can impact MeanSparse’s effectiveness.
VI Conclusions
Adversarial training, a widely accepted defense against evasion attacks, attenuates non-robust feature extractors in a model. Although adversarial training (AT) is effective, this paper demonstrates that there is still an easily accessible capacity for attackers to exploit. We introduce the MeanSparse technique, which partially blocks this capacity in the model available to attackers and enhances adversarial robustness. This technique, easily integrable after training, improves robustness without compromising clean accuracy. By integrating this technique into trained models, we achieve a new record in adversarial robustness in terms of AutoAttack accuracy on CIFAR-10 for both $\ell_{\infty}$ and $\ell_{2}$ untargeted attacks, as well as on CIFAR-100 and ImageNet for $\ell_{\infty}$ untargeted attacks.